okay so it says solve linear system using substitution
4x + 3y = -5
x = y - 3

Answer by stanbon(75887)

using substitution
4x + 3y = -5
x = y - 3
------
Substitute for "x" and solve for "y":
4(y-3) + 3y = -5
4y-12+3y = -5
7y = 7
y = 1
----
Solve for "x":
x = y-3
x = 1-3
x = -2
--------
Cheers,
Stan H.
============